List the on real solutions of the equation below
2x^2-12x+20=0

Answer by stanbon(75887)

2x^2-12x+20=0
-----
x^2 - 6x + 10 = 0
------
Use the Quadratic Formula to get:
x = [6 +- sqrt(36-4*10)]/2
--------
x = [6 +- sqrt(-4)]/2
-------
x = [6 +- 2i]/2
-------
x = 3+i or x = 3-i
----------------
Cheers,
Stan H.
------------